47306 ‘The Sapper’ – In overhaul.
D3452 – Stored awaiting overhaul.
P403D ‘Denise’, 08444, 33110, 37142 & 50042 ‘Triumph’ – In Traffic.
Operational dates this period
28th & 29th Oct Half Price Weekend 37142 running top & tail with Steam.
4th Nov 50042 Lilac TT, celebrating 50 years of 50’s.
2nd Dec 33110 relief loco for Santa specials.
Future running dates
17th & 18th March, Half Price weekend loco TBD top & tail with steam.
8th April Yellow TT, loco TBD x2 round trips
21st April Lilac TT, loco TBD x6 round trips
28th April Lilac TT, loco TBD x6 round trips
29th April Yellow TT, loco TBD x2 round trips
Maintenance for Service Duties
33110 – Serviced.
37142 – Ran the October half term, half price weekend, then laid up for winter. The air relay valve that has been sticking for bogie 1 has been serviced.
50042 – Ran for the 50 years of 50’s in early Nov, a split vacuum hose was replaced on first service run, causing a 15 min delay. 50042 over the winter is in Parkway shed.
Whilst in Parkway shed, the second new KV10 module has been fitted to control the electric train heat generator, also all the small air receivers (Tanks) have been drained with a couple of manual valves now fitted to aid in future draining of these receivers’. Cab 1 desk has been stripped down to allow a drivers wiper control valve exchange to take place.
Daniel Tobbell making and fitting new protection covers for the new KV10 module prior to installation into 50042.
47306 ‘The Sapper’
Work has continued on cleaning, painting of the wheels, bogies, running gear and underframes etc with large parts now in black. A side continues to receive attention with more touching in work, whilst B side is catching up with the filling and sanding etc.
Below each of the cabs the large quantities of expanding foam used to draft proof the loco cabs years ago, is proving quite a task to now remove. This has become necessary to help prevent the corrosion that has taken place ‘hidden’ over the years. Cab 2 internally has seen a lot of effort these last couple of months and is now beginning to look like it could be a cab again, with parts of the desk now primed. New checker plate in places, internal door cover plates renewed and partially painted out.
All of 47306’s window seals have been renewed which has stopped the water entering Cab 2 (Still out in the open, as opposed to Cab 1 which is under shed cover) and this will help the paint out of this cab.
Work has commenced on servicing the machinery with both compressors being cleaned, drained, new oil and compressor one tested on batteries. The air system has been pressurized up to working pressure (144 psi) with a couple of leaks fixed and the triple pump (pressurizes, fuel, lube oil and coolant water) run to first fix leaks / check pressures etc successfully. The external light have all been reinstalled, wired up and tested.